A 1.5 miles long conductor is made up of annealed copper whose diameter is 0.1 inch. Determine the resistance at 30 °C.


1
Expert's answer
2021-03-30T19:54:52-0400

1.5 mil is 2414 m, 0.1 inch is 0.00254 m.

Resistivity of annealed copper at 30°C is 1.68·10-8 Ohm·m.

The resistance is


"R=\frac{\rho L}{A}=\frac{4\rho L}{\pi d^2}=\frac{4(1.68·10^{−8})(2414)}{\pi·0.00254^2}=8\space\Omega."